# Existence results for Leibenson’s equation on Riemannian manifolds

### v1

@@ Abstract
has a weak solution for any $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$.
@@ 1 Introduction
Assume that $pq\geq 1$ and that $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$ is non-negative. Then there exists a non-negative bounded solution $u$ of the Cauchy problem (1.2).
@@
In Section 2, we define the notion of a weak solution of the Leibenson equation (1.1).
@@ 2.1 Definition and basic properties
Note that a priori we do not assume that $u$ is non-negative. In the case when $u$ is not non-negative, we define $u^{q}:=|u|^{q-1}u$. However, since we assume that $u_{0}$ is non-negative, this will imply that $u$ is non-negative (cf. Corollary 3.15).
@@ 3 Estimates of the auxiliary problem
We will show that
$w\in C\left([0,T];L^{2}(\Omega)\right)\text{\ and\ }w\in L^{p}\left((0,T);W_{0}^{1,p}(\Omega)\right)$
and for all $t_{1},t_{2}\in[0,T]$ with $t_{1}<t_{2}$,
@@
where
$w^{h}(\cdot,t)=\left\{\begin{array}[]{ll}\frac{1}{h}\int_{t}^{t+h}w(\cdot,s)ds,&\text{if}~0<t<T-h,\\ 0,&\text{if}~t\geq T-h.\end{array}\right$. (3.37)
Testing with $\psi=w^{h}$ in (3.36), we obtain
@@
Lemma 3.4.
Let $v$ be a weak subsolution of (3.22) and $u$ be a weak supersolution of (3.22). Then, for any $t\in[0,T]$,
$\int_{\Omega}{\left(v-u\right)_{+}^{2}}(t)\leq\int_{\Omega}{\left(v_{0}-u_{0}\right)_{+}^{2}}.$ (3.39)
Proof. We have, for all $0<h<T$ and all $\tau\in[0,T]$,
$\int_{0}^{\tau}\int_{\Omega}{(\partial_{t}v^{h})\psi+\langle[A(v,\nabla v)]^{h},\nabla\psi\rangle}\leq 0,$
and $\psi\in L^{p}\left((0,\tau);W_{0}^{1,p}(\Omega)\right)\cap L^{2}\left(\Omega\times(0,\tau)\right)$, where $v^{h}$ is defined as in (3.37) and the opposite inequality holds for $u$. Applying this and the opposite identity for $u$ with test function $\psi=\left(v^{h}-u^{h}\right)_{+}$, we get
$\int_{0}^{\tau}\int_{\Omega}{(\partial_{t}v^{h}-\partial_{t}u^{h})\psi}+\langle[A(v,\nabla v)]^{h}-[A(u,\nabla u)]^{h},\nabla\psi\rangle\leq 0.$
We obtain
$\int_{0}^{\tau}\int_{\Omega}{\partial_{t}\psi\psi+\langle[A(v,\nabla v)]^{h}-[A(v,\nabla u)]^{h},\nabla\psi\rangle}\leq 0.$ (3.40)
Since
$\int_{0}^{\tau}\int_{\Omega}{\partial_{t}\psi\psi}=\frac{1}{2}\left[\int_{\Omega}{\psi^{2}}\right]_{0}^{\tau},$
we get
$\frac{1}{2}\left[\int_{\Omega}{\psi^{2}}\right]_{0}^{\tau}+\int_{0}^{\tau}{\int_{\Omega}{\langle[A(v,\nabla v)]^{h}-[A(u,\nabla u)]^{h},\nabla\psi\rangle}}\leq 0.$
Sending $h\to 0$ we therefore deduce that
$\frac{1}{2}\left[\int_{\Omega}{\left(v-u\right)_{+}^{2}}\right]_{0}^{\tau}+\int_{0}^{\tau}{\int_{\Omega}{\langle A(v,\nabla v)-A(u,\nabla u),\nabla\left(v-u\right)_{+}\rangle}}\leq 0.$
From the monotonicity property (3.26), we deduce
$\int_{0}^{\tau}{\int_{\Omega}{\langle A(v,\nabla v)-A(u,\nabla u),\nabla\left(v-u\right)_{+}\rangle}}\geq 0,$
and thus,
$\int_{\Omega}{\left(v-u\right)_{+}^{2}}(\tau)\leq\int_{B}{\left(v_{0}-u_{0}\right)_{+}^{2}},$
which proves (3.39).
@@
Assume also that $u_{0}\geq v_{0}$ a.e. in $\Omega$. Then $u\geq v$ a.e. in $\Omega\times(0,T)$.
Proof. By (3.39), we have for any $t\in[0,T]$,
$\int_{\Omega}{\left(v-u\right)_{+}^{2}}(t)\leq\int_{\Omega}{\left(v_{0}-u_{0}\right)_{+}^{2}}.$
Since $u_{0}\geq v_{0}$ a.e. in $\Omega$, we therefore obtain that $u\geq v$ a.e. in $\Omega\times(0,T)$.
Corollary 3.6.
Let $u$ be a weak supersolution of (3.22) and assume that $u_{0}\geq 0$. Then $u\geq\frac{1}{N}$ a.e. in $\Omega\times(0,T)$.
Proof. Applying Corollary 3.5 with subsolution $v=\frac{1}{N}$ and $v_{0}=0$, we get $u\geq v=\frac{1}{N}$ a.e. in $\Omega\times(0,T)$.

### v2

@@ Abstract
has a unique weak solution for any $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$.
@@ 1 Introduction
Assume that $pq\geq 1$ and that $u_{0}\in L^{1}(M)\cap L^{\infty}(M)$ is non-negative. Then there exists a unique non-negative bounded solution $u$ of the Cauchy problem (1.2).
@@
In Section 2, we define the notion of a weak solution of the Leibenson equation (1.1) and prove a comparison principle following a method from [2, 29], which later yields the uniqueness.
@@ 2.1 Definition and basic properties
Note that a priori we do not assume that $u$ is non-negative. In the case when $u$ is not non-negative, we define $u^{q}:=|u|^{q-1}u$. However, since we assume that $u_{0}$ is non-negative, this will imply that $u$ is non-negative (cf. Lemma 2.5).
@@
2.2 Comparison principle
Lemma 2.5.
Let $u$ and $v$ be two weak solutions of (1.1) in $M\times(0,T)$ with inital functions $u_{0},v_{0}$ respectively. Assume that
$u_{0}\leq v_{0}\quad\textnormal{a.e. in}~M.$ (2.22)
Then $u\leq v$ a.e. in $M\times(0,T)$.
The proof of Lemma 2.5 follows from an appropriate choice of test function in the weak formulation (2.15). For that, let $\eta\in C^{\infty}(\mathbb{R})$ be a convex function satisfying
$\eta(z):=\begin{cases}0&\text{when }z\leq 0,\\ z-\frac{1}{2}&\text{when }z\geq 1.\end{cases}$
Note that the derivative $\eta^{\prime}$ approximates the Heaviside function and converges to the sign-function. For $\sigma>0$, let us also define
$\eta_{\sigma,+}(z):=\sigma\eta\!\left(\frac{z}{\sigma}\right),\quad\eta_{\sigma,-}(z):=\sigma\eta\!\left(-\frac{z}{\sigma}\right),$
and
$\varphi_{\sigma,\pm}(z,w):=\int_{w}^{z}\eta^{\prime}_{\sigma,\pm}(s^{q}-w^{q})\,ds,$ (2.23)
where $\eta^{\prime}_{\sigma,\pm}$ is the derivative of $\eta_{\sigma,\pm}$.
Lemma 2.6.
For all $z_{1},z_{2},w\in\mathbb{R}$ and $\sigma>0$ we have
$\varphi_{\sigma,\pm}(z_{1},w)-\varphi_{\sigma,\pm}(z_{2},w)\geq\eta^{\prime}_{\sigma,\pm}(z_{2}^{q}-w^{q})(z_{1}-z_{2}).$ (2.24)
Proof. Since $\eta$ is a smooth, convex function, the same is true for $\eta_{\sigma,\pm}$ and thus $\eta_{\sigma,\pm}^{\prime}$ is non-decreasing. Since $s\mapsto s^{q}$ is also non-decreasing, we see that $s\mapsto\eta_{\sigma,\pm}^{\prime}(s^{q})$ is non-decreasing. Therefore, we have in the case $z_{1}\geq z_{2}$
$\varphi_{\sigma,+}(z_{1},w)-\varphi_{\sigma,+}(z_{2},w)=\int_{z_{2}}^{z_{1}}\eta^{\prime}_{\sigma,+}(s^{q}-w^{q})ds\geq\eta^{\prime}_{\sigma,+}(z_{2}^{q}-w^{q})(z_{1}-z_{2}).$
If $z_{1}<z_{2}$, the same estimate holds as
$\varphi_{\sigma,+}(z_{1},w)-\varphi_{\sigma,+}(z_{2},w)=-\int_{z_{1}}^{z_{2}}\eta^{\prime}_{\sigma,+}(s^{q}-w^{q})\,ds\geq\eta^{\prime}_{\sigma,+}(z_{2}^{q}-w^{q})(z_{1}-z_{2}).$
The proof for $\varphi_{\sigma,-}$ follows by similar arguments so that we conclude (2.24).
Let us also fix a function $\phi\in C^{\infty}_{0}(\mathbb{R})$ satisfying
$\begin{cases}\phi(z)\geq 0\text{ for all }z\in\mathbb{R},\\ \phi(z)=0\text{ for all }|z|\geq 1,\\ \int_{\mathbb{R}}\phi\,dt=1.\end{cases}$
For $\epsilon>0$ and a non-negative function $\gamma\in C^{\infty}_{0}(0,T)$ we define $\gamma_{\epsilon}\in C^{\infty}_{0}(\mathbb{R}^{2})$ by
$\gamma_{\epsilon}(t,s):=\frac{1}{\epsilon}\phi\!\left(\frac{t-s}{\epsilon}\right)\gamma\!\left(\frac{t+s}{2}\right).$ (2.25)
Later we will send $\epsilon$ to 0.
In the following we sometimes omit for simplicity the spatial variable $x$ and write $u(t)=u(x,t)$.
Fix some precompact open subset $\Omega$ of $M$.
Lemma 2.7.
Let $u$ and $v$ be two weak solutions of (2.12) with inital functions $u_{0},v_{0}$ respectively. Let $\omega\in C_{0}^{\infty}(\Omega)$ be non-negative. Then, for all $\epsilon>0$ small enough, all $\sigma>0$, and almost every $s\in(0,T)$, we have
$\int_{0}^{T}\int_{\Omega}\Big($ $-\varphi_{\sigma,+}(u(t),v(s))\partial_{t}\gamma_{\epsilon}(t,s)\omega$ (2.26)
$+\gamma_{\epsilon}(t,s)\eta^{\prime\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla u^{q}(t)-\nabla v^{q}(s)\rangle\omega$
$+\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla\omega\rangle\Big)\,dxdt\leq 0.$
Similarly for small enough $\epsilon$, all $\sigma>0$, and almost every $t\in(0,T)$, it holds that
$\int_{0}^{T}\int_{\Omega}\Big($ $-\varphi_{\sigma,-}(v(s),u(t))\partial_{s}\gamma_{\epsilon}(t,s)\omega$ (2.27)
$-\gamma_{\epsilon}(t,s)\eta^{\prime\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla v^{q}(s)|^{p-2}\nabla v^{q}(s),\nabla u^{q}(t)-\nabla v^{q}(s)\rangle\omega$
$+\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla v^{q}(t)|^{p-2}\nabla v^{q}(t),\nabla\omega\rangle\Big)\,dxds\leq 0.$
Proof. Note that for small enough $\epsilon>0$, only depending on the support of $\gamma$, it holds that
$t\mapsto\gamma_{\epsilon}(t,s)\in C^{\infty}_{0}(0,T)$ (2.28)
for all $s>0$. Also observe that
$\psi(x,t):=\eta^{\prime}_{\sigma,+}(u^{q}(x,t)-v^{q}(x,s))\gamma_{\epsilon}(t,s)\omega(x)\in L^{p}(0,T;W_{0}^{1,p}(\Omega))$
for almost every $s\in(0,T)$.
For notational convenience, we extend $\psi(\cdot,t)=0$ for $t>T$ and $t<0$. Let us consider the Steklov averages
$\psi^{h}(x,t):=\frac{1}{h}\int_{t}^{t+h}\psi(x,\tau)\,d\tau.$ (2.29)
Then it follows from [6] that $\psi^{h}\in L^{p}(0,T;W^{1,p}_{0}(\Omega))$ for all small enough $h>0$ and that $\psi^{h}$ has the time derivative
$\partial_{t}\psi^{h}(t)=\frac{\psi(t+h)-\psi(t)}{h}\in L^{\infty}(M\times(0,T)).$
Hence, we can use $\psi^{h}$ as a test function in (2.15) and obtain
$\int_{0}^{T}\int_{\Omega}u(t)\partial_{t}\psi^{h}(t)\,dxdt=\int_{0}^{T}\int_{\Omega}\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla\psi^{h}(t)\rangle\,dxdt,$
where we used that $\psi_{h}(0)=\psi_{h}(T)=0$ for small enough $h>0$ by (2.28).
By a change of variables and the notational conventions concerning $\psi$ we obtain
$\int_{0}^{T}\int_{\Omega}u(t)\partial_{t}\psi^{h}(t)\,dxdt$ $=\int_{0}^{T}\int_{\Omega}u(t)\frac{\psi(t+h)-\psi(t)}{h}\,dxdt$
$=\int_{0}^{T}\int_{\Omega}-\frac{u(t)-u(t-h)}{h}\psi(t)\,dxdt.$
By the definition of $\psi$ and using (2.24), we further deduce
$\int_{0}^{T}\int_{\Omega}\frac{u(t)-u(t-h)}{h}\psi(t)\,dxdt$ $=\int_{0}^{T}\int_{\Omega}\frac{u(t)-u(t-h)}{h}\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\gamma_{\epsilon}(t,s)\omega\,dxdt$
$\geq\int_{0}^{T}\int_{\Omega}\frac{\varphi_{\sigma,+}(u(t),v(s))-\varphi_{\sigma,+}(u(t-h),v(s))}{h}\gamma_{\epsilon}(t,s)\omega\,dxdt$
$=\int_{0}^{T}\int_{\Omega}-\varphi_{\sigma,+}(u(t),v(s))\frac{\gamma_{\epsilon}(t+h,s)-\gamma_{\epsilon}(t,s)}{h}\omega\,dxdt,$
where we used in the last equality that $v(s)$ does not depend on $t$. Combining this with the computations above we get
$\int_{0}^{T}\int_{\Omega}-\varphi_{\sigma,+}(u(t),v(s))\frac{\gamma_{\epsilon}(t+h,s)-\gamma_{\epsilon}(t,s)}{h}\omega\,dxdt\leq-\int_{0}^{T}\int_{\Omega}\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla\psi^{h}(t)\rangle\,dxdt.$
Note that
$\nabla\psi(t)=\gamma_{\epsilon}(t,s)\eta^{\prime\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))(\nabla u^{q}(t)-\nabla v^{q}(s))\omega+\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\nabla\omega.$
By letting $h\to 0^{+}$ and using the convergence properties of the Steklov averages [6], it follows that
$0$ $\geq\int_{0}^{T}\int_{\Omega}-\varphi_{\sigma,+}(u(t),v(s))\partial_{t}\gamma_{\epsilon}(t,s)\omega+\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla\psi(t)\rangle\,dxdt$
$=\int_{0}^{T}\int_{\Omega}-\varphi_{\sigma,+}(u(t),v(s))\partial_{t}\gamma_{\epsilon}(t,s)\omega$
$\qquad+\gamma_{\epsilon}(t,s)\eta^{\prime\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla u^{q}(t)-\nabla v^{q}(s)\rangle\omega\,$
$\qquad+\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))\langle|\nabla u^{q}(t)|^{p-2}\nabla u^{q}(t),\nabla\omega\rangle dxdt,$
which proves (2.26).
The proof of (2.27), follows from similar arguments by considering the test function
$\psi(x,s)=\eta_{\sigma,-}^{\prime}(v^{q}(s)-u^{q}(t))\gamma_{\epsilon}(t,s)\omega\in L^{p}(0,T;W_{0}^{1,p}(\Omega))$
for small enough $\epsilon>0$, all $\sigma>0$ and almost every $t\in(0,T)$.
Lemma 2.8.
Let $p>1$. Then, for any $\xi,\eta\in T_{x}M$,
In the next lemma we obtain estimates by sending $\sigma\to 0+$ in (2.26) and (2.27).
Lemma 2.9.
Let $u$ and $v$ be two weak solutions of (2.12) with inital functions $u_{0},v_{0}$ respectively. Then, for all $\epsilon>0$ small enough,
$\int_{0}^{T}\int_{0}^{T}\int_{\Omega}(u(t)-v(s))_{+}(\partial_{t}+\partial_{s})\gamma_{\epsilon}(t,s)\omega+\gamma_{\epsilon}(t,s)(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dx\,dt\,ds\geq 0.$ (2.31)
Proof. Integrating (2.26) over $s\in(0,T)$ and (2.27) over $t\in(0,T)$, and adding these two integrals, it follows that
$\int_{0}^{T}\int_{0}^{T}\int_{\Omega}$ $-\varphi_{\sigma,+}(u(t),v(s))\partial_{t}\gamma_{\epsilon}(t,s)\omega-\varphi_{\sigma,-}(v(t),u(s))\partial_{s}\gamma_{\epsilon}(t,s)\omega$
$+\gamma_{\epsilon}(t,s)\eta^{\prime\prime}_{\sigma,+}(u(t)-v(s))\mathcal{A}(\nabla u^{q}(t),\nabla v^{q}(s))$
$-\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdtds\leq 0,$
where we used the notation
$\mathcal{A}(\nabla u^{q},\nabla v^{q})=\left\langle|\nabla u^{q}|^{p-2}\nabla u^{q}-|\nabla v^{q}|^{p-2}\nabla v^{q},\nabla u^{q}-\nabla v^{q}\right\rangle.$
Since $\eta^{\prime\prime}_{\sigma,+}\geq 0$, $\gamma_{\epsilon}\geq 0$ and (2.30), it follows that the term on the second row is non-negative. Hence,
$\int_{0}^{T}\int_{0}^{T}\int_{\Omega}$ $\varphi_{\sigma,+}(u(t),v(s))\partial_{t}\gamma_{\epsilon}(t,s)\omega+\varphi_{\sigma,-}(v(t),u(s))\partial_{s}\gamma_{\epsilon}(t,s)\omega$
$+\gamma_{\epsilon}(t,s)\eta^{\prime}_{\sigma,+}(u^{q}(t)-v^{q}(s))(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdtds\geq 0.$ (2.32)
We now want to let $\sigma\to 0+$ in (2.32). First, note that when $\sigma\to 0+$,
$\eta^{\prime}_{\sigma,+}(z)\to\operatorname{sgn}(z):=\begin{cases}0,&z\leq 0,\\ 1,&z>0,\end{cases}$
uniformly in $\mathbb{R}\setminus[-a,a]$ where $a>0$ is fixed. Hence, as $\sigma\to 0+$, we have $\varphi_{\sigma,+}(z,w)\to(z-w)_{+}$ pointwise in $\mathbb{R}^{2}$. Clearly, we also have $|\varphi_{\sigma,+}(z,w)|\leq(z-w)_{+}$ for all $z,w\in\mathbb{R}$ and $\sigma>0.$ Analogously it also holds that $\varphi_{\sigma,-}(z,w)\to(w-z)_{+}$ and $|\varphi_{\sigma,-}(z,w)|\leq(z-w)_{+}$. Finally, (2.31) follows from the dominated convergence theorem by sending $\sigma\to 0+$ in (2.32), which was to be proved
In the next lemma we let $\epsilon\to 0+$ in (2.31).
Lemma 2.10.
Let $u$ and $v$ be two weak solutions of (2.12) with inital functions $u_{0},v_{0}$ respectively. Then
$\int_{0}^{T}\int_{\Omega}(u(t)-v(t))_{+}\gamma^{\prime}(t)\omega+(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdt\geq 0.$ (2.33)
Proof. Let $\epsilon>0$ be small enough so that (2.31) holds and that $\gamma_{\epsilon}\in C^{\infty}_{0}((0,T)\times(0,T))$. From (2.25) we get
$(\partial_{t}+\partial_{s})\gamma_{\epsilon}(t,s)=\frac{1}{\epsilon}\phi\!\left(\frac{t-s}{\epsilon}\right)\gamma^{\prime}\!\left(\frac{t+s}{2}\right).$
Therefore,
$\int_{0}^{T}\int_{0}^{T}\int_{\Omega}$ $(u(t)-v(s))_{+}\frac{1}{\epsilon}\phi\!\left(\frac{t-s}{\epsilon}\right)\gamma^{\prime}\!\left(\frac{t+s}{2}\right)\omega$
$+\gamma_{\epsilon}(t,s)(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdtds\geq 0.$
By the choice of $\phi$ and by letting $\epsilon\to 0+$ we conclude
$\int_{0}^{T}\int_{\Omega}(u(t)-v(t))_{+}\gamma^{\prime}(t)\omega+\gamma(t)(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdt\geq 0,$
which finishes the proof.
Proof of Lemma 2.5. Let us choose $\gamma$ in (2.33) as a cut-off function, approximating the characteristic function of the interval $t_{1},t_{2}]$, where $0<t_{1}<t_{2}<T$. Hence, we obtain in the limit and using that $u\omega$ and $v\omega\in C([0,T],L^{1}(\Omega))$,
$\int_{\Omega}(u(t_{2})-v(t_{2}))_{+}\omega\,dx+\int_{t_{1}}^{t_{2}}\int_{\Omega}(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdt\leq\int_{\Omega}(u(t_{1})-v(t_{1}))_{+}\omega\,dx.$
For $t_{1}\to 0$ we get
$\int_{\Omega}(u(t_{2})-v(t_{2}))_{+}\omega\,dx+\int_{0}^{t_{2}}\int_{\Omega}(|\nabla u^{q}(t)|^{p-1}+|\nabla v^{q}(t)|^{p-1})|\nabla\omega|\,dxdt\leq\int_{\Omega}(u_{0}-v_{0})_{+}\omega\,dx.$
Finally, choosing $\omega$ such that $\omega\to 1$ and $|\nabla\omega|\to 0$ in $\Omega$, we deduce by (2.22) for any $t\in(0,T)$,
$\int_{\Omega}(u(t)-v(t))_{+}\,dx\leq\int_{\Omega}(u_{0}-v_{0})_{+}\leq 0,$ (2.34)
which implies that $u\leq v$ a.e. in $\Omega\times(0,T)$. Since $\Omega$ was arbitrary, the comparison principle is proved.
@@ 3 Estimates of the auxiliary problem
We will show that $w\in C\left([0,T];L^{2}(\Omega)\right)$ and $w\in L^{p}\left((0,T);W_{0}^{1,p}(\Omega)\right)$ and for all $t_{1},t_{2}\in[0,T]$ with $t_{1}<t_{2}$,
@@
where $w^{h}$ is defined by (2). Testing with $\psi=w^{h}$ in (3.48), we obtain
@@
Let $u$ be a weak solution of (3.35) and assume that $u_{0}\geq 0$. Then $u\geq\frac{1}{N}$ a.e. in $\Omega\times(0,T)$.
Proof. Applying (2.34) with subsolution $v=\frac{1}{N}$ and $v_{0}=0$, we get $u\geq v=\frac{1}{N}$ a.e. in $\Omega\times(0,T)$.